# The systemic recoil velocity distribution and the scale height of isolated millisecond pulsar systems: Implications on neutron star retention fractions in star clusters

### v1

@@ Appendix A Selection of the millisecond pulsar sample
PSR $\mu_{\alpha}$ $\mu_{\delta}$ $\varpi$ Ref. $a$$a$$a$References for proper motions and parallaxes: (1) Bangale et al. (2024), (2) Alam et al. (2020), (3) Ding et al. (2023), (4) Shamohammadi et al. (2024), (5) Curyło et al. (2023), (6) Reardon et al. (2024), (7) van der Wateren et al. (2022), (8) Antoniadis et al. (2023b), (9) Guillemot et al. (2016), (10) Fiore et al. (2023), (11) Kramer et al. (2021), (12) Deller et al. (2019), (13) Deller et al. (2012), (14) Jennings et al. (2018), (15) Konacki & Wolszczan (2003), (16) Yan et al. (2013), (17) Swiggum et al. (2017), (18) Reardon et al. (2021), (19) Chatterjee et al. (2009), (20) Arzoumanian et al. (2018), (21) Deller et al. (2018), (22) Geyer et al. (2023), (23) Ding et al. (2024a), (24) Stovall et al. (2019), (25) Alam et al. (2021), and (26) Crowter et al. (2020).
$d$ $z$ $v_{l}$ $v_{b}$ $v_{\perp}$
($\mathrm{mas\leavevmode\nobreak\ yr^{-1}}$) ($\mathrm{mas\leavevmode\nobreak\ yr^{-1}}$) (mas) (kpc) (kpc) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$)
J0023$+$0923 $-11.00(7)$ $-8.8(1)$ $0.6(1)$ (2), (1) $1.9^{+0.5}_{-0.3}$ $-1.5^{+0.3}_{-0.4}$ $-89^{+16}_{-24}$ $-41^{+9}_{-13}$ $98^{+27}_{-19}$
J0030$+$0451 $-6.15(5)$ $0.4(1)$ $3.04(5)$ (3) $0.329(5)$ $-0.278(5)$ $-9.5(1)$ $12.2(2)$ $15.4(2)$
J0125--2327 $37.14(8)$ $10.5(2)$ $0.8(1)$ (5), (4) $1.2^{+0.2}_{-0.1}$ $-1.2^{+0.1}_{-0.2}$ $35^{+7}_{-5}$ $216^{+34}_{-25}$ $219^{+34}_{-26}$
J0437--4715 $121.4420(6)$ $-71.4717(7)$ $6.43(5)$ (6) $0.156(1)$ $-0.1040(8)$ $46.3(4)$ $85.3(7)$ $97.0(8)$
J0610--2100 $9.11(3)$ $16.45(3)$ $0.7(1)$ (7), (3) $1.5^{+0.3}_{-0.2}$ $-0.46^{+0.07}_{-0.09}$ $-72^{+10}_{-13}$ $95^{+19}_{-14}$ $119^{+24}_{-17}$
J0613--0200 $1.837(2)$ $-10.359(6)$ $1.00(5)$ (8) $1.01(5)$ $-0.163^{+0.008}_{-0.009}$ $44(3)$ $-17.7^{+0.8}_{-0.9}$ $47(3)$
J0614--3329 $0.58(9)$ $-1.9(1)$ $1.5(4)$ (4), (9) $1.0^{+0.7}_{-0.3}$ $-0.4^{+0.1}_{-0.3}$ $22^{+17}_{-8}$ $-8^{+2}_{-3}$ $23^{+18}_{-8}$
J0621$+$1002 $3.27(9)$ $-1.1(3)$ $0.7(1)$ (3) $1.6^{+0.5}_{-0.3}$ $-0.06^{+0.01}_{-0.02}$ $10^{+6}_{-4}$ $17^{+5}_{-4}$ $20^{+8}_{-5}$
J0636$+$5128 $1.1(4)$ $-4.4(7)$ $1.4(3)$ (10), (2) $0.9^{+0.3}_{-0.2}$ $0.28^{+0.11}_{-0.06}$ $8^{+8}_{-5}$ $-5^{+2}_{-3}$ $10^{+8}_{-4}$
J0636--3044 $26.01(9)$ $-19.3(1)$ $4.3(7)$ (4) $0.26^{+0.07}_{-0.04}$ $-0.07^{+0.01}_{-0.02}$ $34^{+10}_{-6}$ $17^{+5}_{-3}$ $39^{+11}_{-7}$
J0645$+$5158 $-7.44(4)$ $1.48(2)$ $0.8(2)$ (2) $1.4^{+0.4}_{-0.3}$ $0.48^{+0.15}_{-0.10}$ $-35^{+5}_{-8}$ $-47^{+9}_{-14}$ $59^{+16}_{-10}$
J0737--3039A $-2.57(3)$ $2.08(4)$ $1.3(1)$ (11) $0.80^{+0.09}_{-0.08}$ $-0.063^{+0.006}_{-0.007}$ $1.4^{+0.6}_{-0.5}$ $-6.2^{+0.5}_{-0.7}$ $6.4^{+0.8}_{-0.6}$
J0751$+$1807 $-2.70(1)$ $-13.27(7)$ $0.85(4)$ (8) $1.18^{+0.06}_{-0.05}$ $0.43(2)$ $55(3)$ $-37(2)$ $67^{+4}_{-3}$
J1012$+$5307 $2.624(3)$ $-25.487(4)$ $1.14(4)$ (3), (8) $0.88(3)$ $0.68(2)$ $82(3)$ $49(2)$ $95(4)$
J1022$+$1001 $-14.92(5)$ $5.61(4)$ $1.39(4)$ (12) $0.72(2)$ $0.56(2)$ $-42(1)$ $-21.8^{+0.8}_{-0.9}$ $47(1)$
J1023$+$0038 $4.76(3)$ $-17.34(4)$ $0.73(3)$ (13) $1.37^{+0.06}_{-0.05}$ $0.98(4)$ $122(5)$ $-33(2)$ $126^{+6}_{-5}$
J1024--0719 $-35.277(5)$ $-48.23(1)$ $0.93(5)$ (3), (8) $1.08(6)$ $0.70(4)$ $61^{+4}_{-3}$ $-291^{+15}_{-18}$ $297^{+17}_{-15}$
J1125--6014 $11.10(4)$ $-12.99(4)$ $1.2(3)$ (4) $2^{+5}_{-1}$ $0.03^{+0.08}_{-0.02}$ $214^{+498}_{-112}$ $-89^{+47}_{-209}$ $231^{+536}_{-121}$
J1227--4853 $-18.8(1)$ $7.30(9)$ $0.5(1)$ (14) $2.9^{+1.5}_{-0.8}$ $0.7^{+0.4}_{-0.2}$ $-183^{+50}_{-87}$ $71^{+37}_{-20}$ $196^{+97}_{-53}$
J1300$+$1240 $45.50(5)$ $-84.70(7)$ $1.41(8)$ (16), (15) $0.72(4)$ $0.69(4)$ $125^{+7}_{-6}$ $-300^{+17}_{-19}$ $325^{+20}_{-18}$
J1400--1431 $17(2)$ $-55(6)$ $4(1)$ (17) $0.7^{+1.2}_{-0.4}$ $0.5^{+0.9}_{-0.3}$ $-2^{+9}_{-23}$ $-199^{+104}_{-344}$ $204^{+351}_{-110}$
J1417--4402 $-4.76(4)$ $-5.10(5)$ $0.20(5)$ (14) $5.3^{+1.6}_{-1.0}$ $1.5^{+0.4}_{-0.3}$ $-14^{+18}_{-9}$ $-100^{+18}_{-23}$ $101^{+22}_{-15}$
J1431--4715 $-11.8(1)$ $-14.5(2)$ $0.5(1)$ (14) $2.6^{+1.4}_{-0.7}$ $0.5^{+0.3}_{-0.1}$ $-148^{+38}_{-67}$ $-114^{+30}_{-61}$ $189^{+95}_{-49}$
J1446--4701 $-4.23(10)$ $-2.4(2)$ $0.7(2)$ (4) $3^{+3}_{-1}$ $0.5^{+0.5}_{-0.2}$ $-7^{+23}_{-3}$ $-11^{+5}_{-11}$ $15^{+13}_{-6}$
J1455--3330 $7.85(1)$ $-1.98(4)$ $1.3(1)$ (8) $0.79^{+0.07}_{-0.06}$ $0.30^{+0.03}_{-0.02}$ $36(2)$ $-23(2)$ $43(3)$
J1518$+$4904 $-0.68(3)$ $-8.53(4)$ $1.24(4)$ (3) $0.81^{+0.03}_{-0.02}$ $0.66(2)$ $-13.4(5)$ $8.6(6)$ $16.0^{+0.8}_{-0.7}$
J1537$+$1155 $1.484(7)$ $-25.29(1)$ $1.07(7)$ (3) $0.95^{+0.07}_{-0.06}$ $0.71^{+0.05}_{-0.04}$ $-85^{+6}_{-7}$ $-56^{+3}_{-4}$ $102^{+8}_{-7}$
J1545--4550 $-0.50(6)$ $2.5(1)$ $0.4(1)$ (18) $5^{+4}_{-2}$ $0.6^{+0.4}_{-0.2}$ $112^{+199}_{-60}$ $40^{+35}_{-16}$ $119^{+204}_{-61}$
J1600--3053 $-0.943(3)$ $-6.92(1)$ $0.72(2)$ (8) $1.39(4)$ $0.39(1)$ $-22.5^{+0.8}_{-0.9}$ $-31.8^{+0.9}_{-1.0}$ $39(1)$
J1614--2230 $4.4(5)$ $-29(3)$ $1.5(1)$ (2) $0.66^{+0.05}_{-0.04}$ $0.23^{+0.02}_{-0.01}$ $-47(8)$ $-72(8)$ $86^{+11}_{-10}$
J1629--6902 $-6.50(3)$ $-8.44(4)$ $0.9(2)$ (4) $1.5^{+0.8}_{-0.4}$ $-0.36^{+0.09}_{-0.20}$ $-46^{+13}_{-25}$ $-1.6(3)$ $46^{+26}_{-13}$
J1640$+$2224 $2.102(4)$ $-11.333(7)$ $0.73(6)$ (3), (8) $1.4(1)$ $0.86^{+0.08}_{-0.07}$ $-43(4)$ $-30(2)$ $53^{+5}_{-4}$
J1643--1224 $6.1(4)$ $3(2)$ $(1)1$ (3) $0.94^{+0.10}_{-0.08}$ $0.34^{+0.04}_{-0.03}$ $38^{+8}_{-7}$ $-13(5)$ $40^{+6}_{-5}$
J1658--5324 $0.0(1)$ $2.9(2)$ $1.3(3)$ (4) $1.4^{+3.2}_{-0.5}$ $-0.16^{+0.06}_{-0.37}$ $33^{+98}_{-9}$ $13^{+33}_{-5}$ $35^{+105}_{-11}$
J1713$+$0747 $4.9215(8)$ $-3.920(2)$ $0.95(6)$ (8), (19) $1.07^{+0.07}_{-0.06}$ $0.45(3)$ $9.8(1)$ $-28(2)$ $29^{+2}_{-1}$
J1723--2837 $-11.73(4)$ $-24.05(3)$ $1.07(4)$ (14) $0.94(4)$ $0.070(3)$ $-109(5)$ $-17.0^{+0.7}_{-0.8}$ $110(5)$
J1730--2304 $20.1(1)$ $-4.7(6)$ $2.0(1)$ (3) $0.51^{+0.03}_{-0.02}$ $0.053(3)$ $27(1)$ $-47^{+2}_{-3}$ $54^{+3}_{-2}$
J1738$+$0333 $7.09(1)$ $5.07(3)$ $0.59(5)$ (3), (8) $1.7^{+0.2}_{-0.1}$ $0.53^{+0.05}_{-0.04}$ $87^{+8}_{-7}$ $-27(2)$ $91^{+8}_{-7}$
J1741$+$1351 $-8.98(2)$ $-7.41(2)$ $0.6(1)$ (20) $1.8^{+0.4}_{-0.3}$ $0.7^{+0.2}_{-0.1}$ $-59^{+10}_{-13}$ $53^{+13}_{-9}$ $79^{+19}_{-13}$
J1744--1134 $18.806(2)$ $-9.39(1)$ $2.58(3)$ (8) $0.388(5)$ $0.0618(7)$ $13.43(4)$ $-38.5(4)$ $40.8(4)$
J1757--5322 $-2.48(8)$ $-10.0(2)$ $1.2(3)$ (4) $1.4^{+1.9}_{-0.5}$ $-0.3^{+0.1}_{-0.4}$ $-52^{+22}_{-67}$ $-13^{+4}_{-13}$ $53^{+71}_{-22}$
J1804--2717 $2.46(2)$ $-16.9(4)$ $1.1(3)$ (8) $6^{+7}_{-4}$ $-0.3^{+0.2}_{-0.3}$ $-252^{+148}_{-94}$ $-284^{+195}_{-347}$ $412^{+310}_{-278}$
J1832--0836 $-8.06(5)$ $-21.0(2)$ $0.5(1)$ (2), (18) $7^{+5}_{-4}$ $0.03(2)$ $-546^{+229}_{-320}$ $-84^{+43}_{-62}$ $551^{+304}_{-232}$
J1853$+$1303 $-1.62(2)$ $-2.96(4)$ $0.53(7)$ (3) $2.0^{+0.4}_{-0.3}$ $0.19^{+0.03}_{-0.02}$ $13^{+3}_{-2}$ $3.5^{+0.8}_{-0.6}$ $14^{+3}_{-2}$
J1857$+$0943 $-2.670(3)$ $-5.428(6)$ $0.89(6)$ (8) $1.15^{+0.09}_{-0.07}$ $0.061^{+0.005}_{-0.004}$ $-7.7^{+0.8}_{-0.9}$ $0.11(4)$ $7.7(9)$
J1909--3744 $-9.523(1)$ $-35.780(5)$ $0.94(2)$ (8) $1.07(2)$ $-0.357^{+0.007}_{-0.008}$ $-176(4)$ $-15.5(3)$ $177(4)$
J1910$+$1256 $0.24(2)$ $-7.10(3)$ $0.36(6)$ (3), (8) $3.2^{+0.9}_{-0.6}$ $0.10^{+0.03}_{-0.02}$ $-12.7^{+4.1}_{-0.9}$ $-52^{+9}_{-14}$ $53^{+13}_{-8}$
J1911$+$1347 $-2.900(5)$ $-3.684(9)$ $0.40(9)$ (8) $3.5^{+2.1}_{-1.0}$ $0.11^{+0.07}_{-0.03}$ $16^{+32}_{-9}$ $16^{+9}_{-4}$ $23^{+31}_{-10}$
J1915$+$1606 $-0.7(1)$ $-0.0(1)$ $0.24(7)$ (21) $6^{+3}_{-2}$ $0.23^{+0.11}_{-0.07}$ $185^{+87}_{-67}$ $20^{+9}_{-6}$ $185^{+86}_{-65}$
J1918--0642 $-7.124(5)$ $-5.96(2)$ $0.71(7)$ (3), (8) $1.5^{+0.2}_{-0.1}$ $-0.23^{+0.02}_{-0.03}$ $-37^{+4}_{-5}$ $23(2)$ $44^{+6}_{-4}$
J1933--6211 $-5.62(1)$ $11.09(3)$ $0.7(2)$ (22), (4) $1.9^{+1.1}_{-0.5}$ $-0.9^{+0.3}_{-0.5}$ $118^{+68}_{-31}$ $68^{+41}_{-20}$ $136^{+80}_{-37}$

### v3

@@ A.1 High-precision samples
PSR solitary? $\mu_{\alpha}$ $\mu_{\delta}$ $\varpi$ Ref. aaaaaaReferences for proper motions and parallaxes: (1) Bangale24, (2) alam20, (3) Ding23, (4) Shamohammadi24, (5) curyo23, (6) reardon24, (7) van-der-Wateren22, (8) Antoniadis23, (9) Guillemot16, (10) fiore23, (11) Kramer21a, (12) Deller19, (13) deller12, (14) Jennings18, (15) Konacki03, (16) Yan13, (17) Swiggum17, (18) reardon21, (19) Chatterjee09, (20) Arzoumanian18a, (21) Deller18, (22) geyer23, (23) ding24c, (24) Stovall19, (25) alam21, and (26) Crowter20.
$d$ $z$ $v_{l}$ $_{b} vv_{\perp}$
(Y/N) ($\mathrm{mas\penalty 10000\ yr^{-1}}$) ($\mathrm{mas\penalty 10000\ yr^{-1}}$) (mas) (kpc) (kpc) ($\mathrm{km\penalty 10000\ s^{-1}}$) ($\mathrm{km\penalty 10000\ s^{-1}}$) ($\mathrm{km\penalty 10000\ s^{-1}}$)
J0023++0923 N 11.00(7)-11.00(7) 8.8(1)-8.8(1) $0.6(1)$ (1), (2) $1.9^{+0.5}_{-0.3}$ $1.5^{+0.3}_{-0.4}$ $88^{+16}_{-24}$ $41^{+9}_{-13}$ $98^{+28}_{-19}$
J0030++0451 Y 6.15(5)-6.15(5) $0.4(1)$ $3.04(5)$ (3) $0.329(5)$ 0.278(5)-0.278(5) 9.5(1)-9.5(1) $12.2(2)$ $15.4(2)$
J0125-2327 N $37.14(8)$ $10.5(2)$ $0.8(1)$ (4), (5) $1.2^{+0.2}_{-0.1}$ $1.2^{+0.1}_{-0.2}$ $35^{+7}_{-5}$ $16^{+34}_{-25} 22$$20^{+34}_{-26}$
J0437-4715 N $121.4420(6)$ 71.4717(7)-71.4717(7) $6.43(5)$ (6) $0.156(1)$ 0.1040(8)-0.1040(8) $46.3(4)$ $85.3(7)$ $97.0(8)$
J0610-2100 N $9.11(3)$ $16.45(3)$ $0.7(1)$ (3), (7) $1.5^{+0.3}_{-0.2}$ $0.46^{+0.07}_{-0.09}$ $72^{+10}_{-14}$ $95^{+20}_{-14}$ $119^{+23}_{-17}$
J0613-0200 N $1.837(2)$ 10.359(6)-10.359(6) $1.00(5)$ (8) $1.01(5)$ $0.163^{+0.008}_{-0.009}$ $44(3)$ $17.7^{+0.8}_{-0.9}$ $47(3)$
J0614-3329 N $0.58(9)$ 1.9(1)-1.9(1) $1.5(4)$ (4), (9) $1.0^{+0.7}_{-0.3}$ $0.4^{+0.1}_{-0.3}$ $21^{+18}_{-8}$ $8^{+1}_{-4}$ $23^{+18}_{-8}$
J0621++1002 N $3.27(9)$ 1.1(3)-1.1(3) $0.7(1)$ (3) $1.6^{+0.5}_{-0.3}$ $0.06^{+0.01}_{-0.02}$ $10^{+6}_{-4}$ $17^{+5}_{-3}$ $20^{+8}_{-5}$
J0636++5128 N $1.1(4)$ 4.4(7)-4.4(7) $1.4(2)$ (2), (10) $0.8^{+0.2}_{-0.1}$ $0.26^{+0.06}_{-0.04}$ $7^{+5}_{-4}$ 5(2)-5(2) $9^{+5}_{-3}$
J0636-3044 Y $26.01(9)$ 19.3(1)-19.3(1) $4.3(7)$ (4) $26^{+0.07}_{-0.04} 0.0$$07^{+0.01}_{-0.02}$ $34^{+10}_{-6}$ $17^{+5}_{-3}$ $38^{+11}_{-7}$
J0645++5158 Y $1.55(2)$ 7.47(4)-7.47(4) $0.8(2)$ (2) $1.4^{+0.4}_{-0.3}$ $0.47^{+0.13}_{-0.09}$ $39^{+14}_{-9}$ 10(2)-10(2) $40^{+14}_{-9}$
J0737-3039A N 2.57(3)-2.57(3) $2.08(4)$ $1.3(1)$ (11) $80^{+0.09}_{-0.08} 0.0$$063^{+0.006}_{-0.007}$ $1.4^{+0.6}_{-0.5}$ $6.2^{+0.5}_{-0.7}$ $6.4^{+0.8}_{-0.6}$
J0751++1807 N 2.70(1)-2.70(1) 13.27(7)-13.27(7) $0.85(4)$ (8) $1.18^{+0.06}_{-0.05}$ $0.43(2)$ $55(3)$ 37(2)-37(2) $67^{+4}_{-3}$
J1012++5307 N $2.624(3)$ 25.487(4)-25.487(4) $1.14(4)$ (8), (3) $0.88(3)$ $0.68(2)$ $82(3)$ $49(2)$ $95(4)$
J1022++1001 N 14.92(5)-14.92(5) $5.61(4)$ $1.39(4)$ (12) $0.72(2)$ $0.56(2)$ $42.2^{+1.0}_{-1.1}$ $21.8^{+0.8}_{-0.9}$ $47(1)$
J1023++0038 N $4.76(3)$ 17.34(4)-17.34(4) $0.73(3)$ (13) $1.37^{+0.06}_{-0.05}$ $0.98(4)$ $122(5)$ 33(2)-33(2) $126^{+6}_{-5}$
J1125-6014 N $11.09(3)$ 13.00(3)-13.00(3) $1.2(3)$ (4) $2^{+5}_{-1}$ $0.03^{+0.08}_{-0.02}$ $210^{+484}_{-108}$ $87^{+45}_{-203}$ $234^{+536}_{-123}$
J1227-4853 N 18.8(1)-18.8(1) $7.30(9)$ $0.5(1)$ (14) $2.9^{+1.5}_{-0.8}$ $0.7^{+0.4}_{-0.2}$ $183^{+50}_{-91}$ $71^{+39}_{-20}$ $195^{+97}_{-52}$
J1300++1240 N $45.50(4)$ 84.70(7)-84.70(7) $1.41(8)$ (16), (15) $0.72(4)$ $0.69(4)$ $125^{+7}_{-6}$ $300^{+17}_{-19}$ $325^{+20}_{-18}$
J1400-1431 N $17(2)$ 55(6)-55(6) $4(1)$ (17) $0.7^{+1.2}_{-0.4}$ $0.5^{+0.9}_{-0.3}$ $2^{+9}_{-22}$ $201^{+105}_{-341}$ $201^{+348}_{-107}$
J1417-4402 N 4.76(4)-4.76(4) 5.10(5)-5.10(5) $0.20(5)$ (14) $5.3^{+1.6}_{-1.0}$ $1.5^{+0.4}_{-0.3}$ $14^{+18}_{-9}$ $100^{+18}_{-23}$ $101^{+22}_{-16}$
J1431-4715 N 11.8(1)-11.8(1) 14.5(2)-14.5(2) $0.5(1)$ (14) $2.6^{+1.4}_{-0.7}$ $0.5^{+0.3}_{-0.1}$ $49^{+38}_{-67} 11$$14^{+31}_{-62}$ $188^{+91}_{-49}$
J1446-4701 N 4.24(9)-4.24(9) 2.4(2)-2.4(2) $0.7(2)$ (4) $3^{+3}_{-1}$ $0.5^{+0.5}_{-0.2}$ $8^{+22}_{-3}$ $12^{+5}_{-10}$ $15^{+14}_{-6}$
J1455-3330 N $7.85(1)$ 1.98(4)-1.98(4) $1.3(1)$ (8) $0.79^{+0.07}_{-0.06}$ $0.30^{+0.03}_{-0.02}$ $36(2)$ 23(2)-23(2) $43(3)$
J1518++4904 N 0.68(3)-0.68(3) 8.53(4)-8.53(4) $1.24(4)$ (3) $0.81^{+0.03}_{-0.02}$ $0.66(2)$ 13.5(5)-13.5(5) $8.6^{+0.6}_{-0.5}$ $16.0^{+0.8}_{-0.7}$
J1537++1155 N $1.484(7)$ 25.29(1)-25.29(1) $1.07(7)$ (3) $0.95^{+0.07}_{-0.06}$ $0.71^{+0.05}_{-0.04}$ $85^{+6}_{-7}$ $56^{+3}_{-4}$ $102^{+8}_{-7}$
J1545-4550 N 0.51(7)-0.51(7) $2.5(1)$ $0.4(1)$ (4), (18) $5^{+4}_{-2}$ $0.6^{+0.4}_{-0.2}$ $112^{+197}_{-60}$ $40^{+34}_{-16}$ $120^{+202}_{-61}$
J1600-3053 N 0.943(3)-0.943(3) 6.92(1)-6.92(1) $0.72(2)$ (8) $1.39(4)$ $0.39(1)$ $22.5^{+0.8}_{-0.9}$ 31.8(9)-31.8(9) $39(1)$
J1614-2230 N $3.9(1)$ 32.1(5)-32.1(5) $1.5(1)$ (2) $0.66^{+0.05}_{-0.04}$ $0.23^{+0.02}_{-0.01}$ $55^{+4}_{-5}$ $78^{+5}_{-6}$ $95(7)$
J1629-6902 Y 6.50(3)-6.50(3) 8.44(4)-8.44(4) $0.9(2)$ (4) $1.5^{+0.8}_{-0.4}$ $0.36^{+0.09}_{-0.20}$ $46^{+13}_{-25}$ 1.6(3)-1.6(3) $45^{+25}_{-12}$
J1640++2224 N $2.102(4)$ 11.333(7)-11.333(7) $0.73(6)$ (8), (3) $1.4(1)$ $0.86^{+0.08}_{-0.07}$ 44(4)-44(4) 30(2)-30(2) $53^{+5}_{-4}$
J1643-1224 N $6.1(4)$ $3(2)$ $1.1(1)$ (3) $0.94^{+0.10}_{-0.08}$ $0.34^{+0.04}_{-0.03}$ $38^{+8}_{-7}$ 13(5)-13(5) $40^{+6}_{-5}$
J1658-5324 Y $0.0(1)$ $2.9(2)$ $1.3(3)$ (4) $1.4^{+3.2}_{-0.5}$ $0.16^{+0.06}_{-0.37}$ $33^{+95}_{-9}$ $13^{+33}_{-5}$ $35^{+105}_{-11}$
J1713++0747 N $4.9215(8)$ 3.920(2)-3.920(2) $0.95(6)$ (8), (19) $1.07^{+0.07}_{-0.06}$ $0.45(3)$ $9.8(1)$ 28(2)-28(2) $29^{+2}_{-1}$
J1723-2837 N 11.73(4)-11.73(4) 24.05(3)-24.05(3) $1.07(4)$ (14) $0.94(4)$ $0.070(3)$ 109(5)-109(5) $7.1^{+0.7}_{-0.8} 11$0(5)110(5)
J1730-2304 Y $20.1(1)$ 4.7(6)-4.7(6) $2.0(1)$ (3) $51^{+0.03}_{-0.02} 0.0$53(3)0.053(3) $27^{+2}_{-1}$ $46^{+2}_{-3}$ $54^{+3}_{-2}$
J1738++0333 N $7.09(1)$ $5.07(3)$ $0.59(5)$ (8), (3) $1.7^{+0.2}_{-0.1}$ $0.53^{+0.05}_{-0.04}$ $87^{+8}_{-7}$ 27(2)-27(2) $91^{+8}_{-7}$
J1741++1351 N 8.97(1)-8.97(1) 7.44(2)-7.44(2) $0.6(1)$ (2), (20) $1.8^{+0.4}_{-0.3}$ $0.7^{+0.2}_{-0.1}$ $59^{+10}_{-13}$ $53^{+13}_{-9}$ $79^{+19}_{-13}$
J1744-1134 Y $18.806(2)$ 9.39(1)-9.39(1) $2.58(3)$ (8) $0.388(5)$ $0.0618(7)$ $13.43(4)$ 38.5(4)-38.5(4) $40.8(4)$
J1757-5322 N 2.48(8)-2.48(8) 10.0(2)-10.0(2) $1.2(3)$ (4) $1.4^{+1.9}_{-0.5}$ $0.3^{+0.1}_{-0.4}$ $53^{+22}_{-67}$ $13^{+4}_{-14}$ $54^{+72}_{-23}$
J1804-2717 N $2.46(2)$ 16.9(4)-16.9(4) $1.1(3)$ (8) $6^{+7}_{-4}$ $0.3^{+0.2}_{-0.3}$ $251^{+148}_{-91}$ $279^{+191}_{-348}$ $412^{+307}_{-276}$
J1832-0836 Y 8.06(5)-8.06(5) 21.0(2)-21.0(2) $0.5(1)$ (2), (18) $6^{+5}_{-3}$ $0.03^{+0.02}_{-0.01}$ $493^{+214}_{-269}$ $71^{+35}_{-62}$ $497^{+289}_{-217}$
J1853++1303 N 1.62(2)-1.62(2) 2.96(4)-2.96(4) $0.53(7)$ (3) $2.0^{+0.4}_{-0.3}$ $0.19^{+0.03}_{-0.02}$ $13^{+3}_{-2}$ $3.5^{+0.8}_{-0.6}$ $14^{+3}_{-2}$
J1857++0943 N 2.670(3)-2.670(3) 5.428(6)-5.428(6) $0.89(6)$ (8) $1.15^{+0.09}_{-0.07}$ $0.061^{+0.005}_{-0.004}$ $7.7^{+0.8}_{-0.9}$ $0.11^{+0.05}_{-0.04}$ $7.7^{+0.9}_{-0.8}$
J1909-3744 N 9.523(1)-9.523(1) 35.780(5)-35.780(5) $0.94(2)$ (8) $1.07(2)$ $0.357^{+0.007}_{-0.008}$ 177(4)-177(4) 15.5(3)-15.5(3) $177(4)$
J1910++1256 N $0.24(2)$ 7.10(3)-7.10(3) $0.36(6)$ (8), (3) $3.2^{+0.9}_{-0.6}$ $0.10^{+0.03}_{-0.02}$ $12.7^{+4.2}_{-0.9}$ $52^{+9}_{-14}$ $54^{+13}_{-8}$
J1911++1347 Y 2.900(5)-2.900(5) 3.684(9)-3.684(9) $0.40(9)$ (8) $3.5^{+2.1}_{-1.0}$ $0.11^{+0.07}_{-0.03}$ $16^{+33}_{-9}$ $16^{+9}_{-4}$ $23^{+32}_{-9}$
J1915++1606 N 0.7(1)-0.7(1) 0.0(1)-0.0(1) $0.24(7)$ (21) $6^{+3}_{-2}$ $0.23^{+0.11}_{-0.07}$ $184^{+86}_{-66}$ $20^{+9}_{-6}$ $183^{+86}_{-64}$
J1918-0642 N 7.124(5)-7.124(5) 5.96(2)-5.96(2) $0.71(7)$ (8), (3) $1.5^{+0.2}_{-0.1}$ $0.23^{+0.02}_{-0.03}$ $37^{+4}_{-5}$ $23^{+3}_{-2}$ $44^{+5}_{-4}$
J1933-6211 N 5.62(1)-5.62(1) $11.09(3)$ $0.7(2)$ (4), (22) $1.9^{+1.1}_{-0.5}$ $0.9^{+0.3}_{-0.5}$ $118^{+68}_{-31}$ $68^{+42}_{-20}$ $136^{+79}_{-37}$
J1939++2134 Y $0.074(2)$ 0.410(3)-0.410(3) $0.35(3)$ (3) $3.0^{+0.3}_{-0.2}$ 0.015(1)-0.015(1) $80^{+9}_{-7}$ $3.8^{+0.3}_{-0.4}$ $80^{+9}_{-8}$